### v1

# Selective Mixup for Debiasing Question Selection in Computerized Adaptive Testing

Abstract.
Computerized Adaptive Testing is a widely used technology for evaluating learners’ proficiency in online education platforms. By leveraging prior estimates of proficiency to select questions and updating the estimates iteratively based on responses, CAT enables personalized learner modeling and has attracted substantial attention. Despite this progress, most existing works focus primarily on improving diagnostic accuracy, while overlooking the selection bias inherent in the adaptive process. Selection Bias arises because the question selection is strongly influenced by the estimated proficiency, such as assigning easier questions to learners with lower proficiency and harder ones to learners with higher proficiency. Since the selection depends on prior estimation, this bias propagates into the diagnosis model, which is further amplified during iterative updates, leading to misalignment and biased predictions. Moreover, the imbalanced nature of learners’ historical interactions often exacerbates the bias in diagnosis models. To address this issue, we propose a debiasing framework consisting of two key modules: Cross-Attribute Examinee Retrieval and Selective Mixup-based Regularization. First, we retrieve balanced examinees with relatively even distributions of correct and incorrect responses and use them as neutral references for biased examinees. Then, mixup is applied between each biased examinee and its matched balanced counterpart under label consistency. This augmentation enriches the diversity of bias-conflicting samples and smooths selection boundaries. Finally, extensive experiments on two benchmark datasets with multiple advanced diagnosis models demonstrate that our method substantially improves both the generalization ability and fairness of question selection in CAT.
Computerized Adaptive Testing, Intelligent Education, Bias Mitigation
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Korea.booktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3761121isbn: 979-8-4007-2040-6/2025/11ccs: Applied computing Computer-assisted instruction
1. Introduction
Figure 1. Training Process of Data-Driven Question Selection Module.
Computerized Adaptive Testing (CAT) enables accurate proficiency estimation with fewer questions by iteratively selecting questions and updating learner states (liu2024survey). While early CAT relied on heuristic strategies such as Maximum Fisher Information (lord2012mfi), recent advances have shifted toward data-driven approaches that learn the Question Selection Module from large-scale interaction data (ghosh2021bobcat; wang2023gmocat; yu2024ucats), achieving strong diagnostic performance. However, research on mitigating bias in CAT remains limited (kwon2023useraif; hort2024biassurvey). Bias arising from data imbalance or adaptive selection may lead to systematically easier or harder questions being assigned, thereby distorting proficiency estimates and undermining fairness. Prior work such as UserAIF (kwon2023useraif) alleviates bias by recalibrating question parameters, but it overlooks the effect of biased training data on the Question Selection Module itself.
The origins and consequences of biased selection remain insufficiently explored, leaving a gap in the development of fair and robust CAT systems. This work addresses the gap by examining how biased selection arises during training and affects both accuracy and fairness, with a focus on mitigating bias within data-driven Question Selection Modules. Recent Question Selection Modules often adopt a bi-level optimization framework (ghosh2021bobcat; yu2024ucats), where each examinee’s interactions are split into a support set and a meta set. The support set is used by the Cognitive Diagnosis Module to estimate proficiency, while the meta set provides feedback for training the Question Selection Module. As illustrated in Figure 1, at each step $t$, the Question Selection Module selects the next question based on the examinee’s history, the CDM updates proficiency after the response, and predictions on the meta set yield feedback to optimize the Question Selection Module. This process repeats for $T$ steps to iteratively refine question selection.
In this process, the meta set guides how the Question Selection Module is trained, aiming to align the distribution of selected questions with that of the meta set to improve diagnostic accuracy. In practice, however, the meta set is inherently imbalanced. As training progresses, the Question Selection Module overfits these skewed patterns, causing a divergence between selected and unbiased distributions, which exacerbates selection bias.
(a) Distribution of the meta set and selected set.
(b) Biased proficiency estimation.
Figure 2. The imbalanced distribution in the meta set can lead to a shift in the selection set, resulting in biased estimates of examinee proficiency.
Examinees differ in their proportion of correct responses: some perform well, others poorly, while some are more balanced. We characterize each examinee by this correct-response ratio. As shown in Figure 2(a), the distribution of meta sets and selected data in the NIPS-EDU dataset varies with this ratio [^1] https://eedi.com/projects/neurips-education-challenge, with density curves estimated by histograms and cubic spline smoothing. For high-ratio examinees, the meta set centers around 0.6 but shifts toward 1.0 after training, indicating a bias toward reinforcing strong performance. For low-ratio examinees, the distribution similarly shifts toward 0.0, amplifying underperformance. Thus, an already imbalanced meta set becomes more skewed during training. This imbalance introduces selection bias: the Question Selection Module tends to favor easier questions for high-ratio examinees and harder ones for low-ratio examinees, regardless of their true proficiency. While data-driven CAT models achieve strong empirical performance, such biased behavior contradicts the core objective of CAT and leads to distorted outcomes. As illustrated in Figure 2(b), this results in unreliable estimates, systematically overestimating high performers and underestimating low performers.
Selection bias in data-driven CAT arises from imbalanced meta sets, where rare cases—such as correct responses from low-performing students or incorrect responses from high-performing students—are underrepresented. This imbalance leads the Question Selection Module to overfit common patterns and hinders generalization. Since CAT jointly optimizes the Cognitive Diagnosis Module and the Question Selection Module, it is difficult to isolate the source of bias. To focus on selection bias, we fix the Cognitive Diagnosis Module and address imbalance in the Question Selection Module by augmenting rare cases with interactions from balanced examinees with matching labels. Through selective mixup of these samples, we construct a more balanced training distribution that guides the Question Selection Module to learn smoother, less biased decision boundaries, thereby improving robustness and fairness under unbiased evaluation.
  • Formally identify and analyze the problem of selection bias in data-driven CAT, tracing its origin to imbalanced meta set distributions during training.
  • Propose a selective cross-attribute mixup strategy to augment underrepresented interactions and enhance the generalization of the Question Selection Module.
  • Validate the effectiveness of the proposed method through extensive experiments on two benchmark datasets, demonstrating strong bias-mitigation capability .
2. Related Work
2.1. CAT and Bias in Cognitive Modeling
Computerized Adaptive Testing (CAT) adaptively selects questions to estimate examinees’ knowledge mastery (chang2015survey; liu2024survey), typically consisting of a Cognitive Diagnosis Module (CDM) to infer knowledge states and a Question Selection Module to determine the next qustion. Early CDMs were built on Item Response Theory  (embretson2013irt) and its multidimensional extension  (ackerman2014mirt), while recent deep learning models such as NCDM (wang2020ncd), KaNCD (wang2022kancd), RCD (gao2021rcd), and DCD (chen2023dcd) have greatly improved representation and prediction. For Question Selection Module, heuristic approaches such as Maximum Fisher Information (lord2012mfi), KL-based selection  (chang1996kli), MAAT (bi2020maat), and BECAT (zhuang2023becat) are simple and interpretable but limited in modeling complex interactions. More recent data-driven methods leverage bi-level optimization or reinforcement learning, including BOBCAT (ghosh2021bobcat), NCAT (zhuang2022ncat), GMOCAT (wang2023gmocat), LACAT (cheng2024lacat), and UATS (yu2024ucats). However, data-driven methods often overlook bias in the data and selection process, which may compromise robustness in CAT.
Bias has been increasingly recognized in CDMs, where predictions may differ systematically across sensitive groups such as gender or socioeconomic status (zhang2024psrcf; zhang2024faircd; xu2025fairwisa). Methods like adversarial learning or causal reasoning have been proposed to mitigate such disparities. However, bias mitigation in CAT remains relatively underexplored, and research on bias mitigation in CAT is still very limited. In practice, CAT interaction data is inherently subject to selection bias. Training data-driven CAT models directly on such biased data can distort parameter estimation and compromise the validity of the system. Prior work such as UserAIF (kwon2023useraif) uses influence functions to identify less biased samples and recalibrate the question parameters in Cognitive Diagnosis Module.
Our distinction. Previous studiy focuses on question-side bias, while our work targets the Question Selection Module, mitigating bias from adaptive selection to improve fairness and robustness in CAT, directly correcting bias at its source rather than applying debiasing after question selection.
2.2. Bias Mitigation Methods
While Machine Learning systems can relieve humans from tedious tasks and perform complex calculations at high speed, they are only as reliable as the data on which they are trained (barocas2016big). Since algorithms are not intentionally designed to incorporate bias, they often replicate or even amplify the spurious correlations present in real-world data, raising concerns about fairness and accountability (pedreshi2008discrimination; van2022overcoming). To address these issues, a variety of strategies have been proposed. Invariant Risk Minimization (IRM) seeks environment-invariant predictors (arjovsky2019irm), while GroupDRO minimizes worst-group risk but requires group labels (sagawa2019groupDRO). Reweighting strategies adjust sample importance to counter imbalance (japkowicz2000reweight), yet they only reshape optimization weights without enriching data diversity. In contrast, Mixup-based methods (zhang2018mixup) regularize models by interpolating samples and labels, with extensions such as Manifold Mixup (verma2019manifold) and Co-Mixup (kim2021co) improving hidden representation learning and pair selection. Building on this line of work, our method adopts a selective mixup strategy to mitigate selection bias in CAT, leveraging the simplicity and effectiveness of Mixup to enhance generalization.
3. Preliminaries
3.1. Computerized Adaptive Testing
CAT reduces the burden on examinees by minimizing the number of questions while ensuring that the estimated proficiency $\boldsymbol{\theta}$ aligns with the true proficiency $\boldsymbol{\theta}_{0}$. It consists of a Cognitive Diagnosis Module, which estimates proficiency from responses, and a Question Selection Module, which adaptively selects the next item. Their feedback loop enables efficient and accurate assessment.
3.1.1. Cognitive Diagnosis Module
The Cognitive Diagnosis Module (CDM) estimates an examinee’s proficiency $\boldsymbol{\theta}$ by fitting their historical responses and predicting performance on unseen questions. Formally, the model outputs the probability $p\in[0,1]$ that an examinee answers a given item correctly. A representative neural network–based CDM is NCDM (wang2020ncd), which leverages neural networks to capture complex interactions between examinees and items. The model incorporates several components: an examinee proficiency vector $\boldsymbol{\theta}\in(0,1)^{1\times K}$, a question concept vector $\boldsymbol{Q}_{i}\in{0,1}^{1\times K}$ indicating the required knowledge concepts, a knowledge difficulty vector $\boldsymbol{h}^{diff}\in(0,1)^{1\times K}$, a scalar discrimination parameter $h^{disc}\in(0,1)$, and other trainable neural network parameters.
Interaction Function.
The first interaction layer is inspired by the Multidimensional question Response Theory (MIRT) (ackerman2014mirt) model, and is defined as:
(1) $\boldsymbol{x}=\boldsymbol{Q}_{i}\circ(\boldsymbol{\theta}-\boldsymbol{h}^{diff})\times h^{disc},$
where $\circ$ denotes the element-wise product.The intermediate layers consist of two fully connected layers followed by an output layer:
(2) $\boldsymbol{f}_{1}$ $=\varphi(\boldsymbol{W}_{1}\boldsymbol{x}^{T}+\boldsymbol{b}_{1}),$
(3) $\boldsymbol{f}_{2}$ $=\varphi(\boldsymbol{W}_{2}\boldsymbol{f}_{1}+\boldsymbol{b}_{2}),$
(4) $p$ $=\varphi(\boldsymbol{W}_{3}\boldsymbol{f}_{2}+\boldsymbol{b}_{3}),$
where $\varphi$ denotes the activation function, typically implemented as a sigmoid. The monotonicity assumption requires that the probability of correctly answering an item increases monotonically with respect to any dimension of a student’s knowledge proficiency. To enforce this property, NCDM adopts a simple yet effective strategy by constraining all elements of $\boldsymbol{W}_{1}$, $\boldsymbol{W}_{2}$, and $\boldsymbol{W}_{3}$ to be positive.
Loss Function.
The cross-entropy loss is adopted between the predicted probability $M(q;\boldsymbol{\theta},\zeta_{q},\zeta_{n})=p\in[0,1]$ and the ground-truth response $y$:
(5) $\ell(y,M(q;\boldsymbol{\theta},\zeta_{q},\zeta_{n}))=-\big(y\log p+(1-y)\log(1-p)\big),$
where $q$ denotes the exercise, $\boldsymbol{\theta}$ represents the examinee proficiency, $\zeta_{q}$ corresponds to the question parameters (e.g., concept vector, difficulty, discrimination), and $\zeta_{n}$ denotes the trainable neural network parameters. After training, the estimated proficiency vector $\boldsymbol{\theta}$ serves as the diagnostic result, representing the examinee’s proficiency over the $K$ knowledge concepts.
3.1.2. Question Selection Module
Existing data-driven Question Selection Module typically employ a bi-level optimization framework (ghosh2021bobcat), where a selection network is trained on large-scale examinee data. The dataset $\mathcal{D}$ is divided into a support set $\mathcal{D}_{u}$ and a meta set $\mathcal{D}_{m}$ used to compute the loss or reward. To illustrate, we briefly introduce BOBCAT (ghosh2021bobcat) as an example. At each step $t$, the selection network $\pi(\cdot)$ chooses a question $q_{t}\in\mathcal{D}_{u}$ based on the examinee’s current state $\boldsymbol{s_{t}}$, which is encoded from their interaction history $r=\{(q_{1},y_{1}),\ldots,(q_{t-1},y_{t-1})\}$, where $y_{i}\in\{0,1\}$ indicates whether the response was correct or incorrect.
The state $\boldsymbol{s_{t}}$ is a vector of length equal to the number of questions, where each entry is in $\{-1,0,1\}$, representing incorrect, unattempted, and correct responses, respectively. Formally, the selection process is defined as:
(6) $q_{t}=\pi(\boldsymbol{s_{t}},\mathcal{D}_{u};\zeta_{s})$
where $\zeta_{s}$ denotes the parameters of the selection network.
To train the selection network, a bi-level optimization framework is adopted. The outer loop focuses exclusively on updating the selection network $\zeta_{s}$, since the Question Selection Module is assumed to be the main source of bias. The Cognitive Diagnosis Module is kept fixed, with all parameters $(\zeta_{q},\zeta_{n})$ pre-trained and frozen, while only the proficiency representation $\boldsymbol{\theta}$ is updated during inner optimization. After training, the selection network generates the next question for each examinee based solely on their current response state. The overall bi-level optimization procedure for data-driven question selection is summarized as follows:
Inner Optimization.
Given a selected question sequence, the examinee’s proficiency $\boldsymbol{\theta}_{i}$ is estimated by minimizing the prediction loss on their responses in $\mathcal{D}_{u}$ with fixed question parameters $(\zeta_{q},\zeta_{n})$, as defined in Eq. (5). Through this optimization, $\boldsymbol{\theta}_{i}$ is updated to obtain the estimated proficiency $\boldsymbol{\theta}_{i}^{*}$:
(7) $\boldsymbol{\theta}_{i}^{*}=\arg\min_{\boldsymbol{\theta}_{i}}\sum_{t=1}^{T}\ell\!\left(y_{i,q_{t}},M\left(q_{t};\boldsymbol{\theta}_{i},\zeta_{q},\zeta_{n}\right)\right),$
where $T$ denotes the time step, i.e., the number of selected questions.
Outer Optimization.
The selection network $\zeta_{s}$ is then updated by minimizing the loss on the meta set $\mathcal{D}_{m}$ with the fixed proficiency $\boldsymbol{\theta}_{i}^{*}$ and pre-trained CDM parameters $(\zeta_{q},\zeta_{n})$:
(8) $\min_{\zeta_{s}}\frac{1}{N}\sum_{i=1}^{N}\sum_{q_{j}\in\mathcal{D}_{m}}\ell\!\left(y_{i,q_{j}},M\left(q_{j};\boldsymbol{\theta}_{i}^{*},\zeta_{q},\zeta_{n}\right)\right),$
where $N$ denotes the number of examinees.
3.2. Notation and Problem Setting
We divide the examinees in $\mathcal{D}_{m}$ into three attributes: $\mathcal{D}_{m}^{\text{A}},\ \mathcal{D}_{m}^{\text{B}},\ \mathcal{D}_{m}^{\text{C}}$, based on their proportion of correct responses. These groups represent different meta set distributions. Let $\mathcal{A}=\{\text{A},\text{B},\text{C}\}$ denote the attribute categories.
  • A: the proportion of correct responses is in the range $[0,0.4]$, where incorrect responses dominate.
  • B: the proportion of correct responses falls within $(0.4,0.6]$, indicating a roughly balanced distribution.
  • C: the proportion of correct responses is in the range $(0.6,1.0]$, where correct responses dominate.
We also define the label set as $\mathcal{Y}=\{0,1\}$, where 0 denotes an incorrect response and $1$ denotes a correct response. A group is defined as a unique combination of an attribute $a\in\mathcal{A}$ and a label $y\in\mathcal{Y}$. Based on this group definition, we further categorize the interactions into three types:
  • Bias-aligned samples: Interactions that conform to the dominant attribute pattern and are typically easy to learn. This includes interactions from $\mathcal{D}_{m}^{\text{A}}$ with label 0, and from $\mathcal{D}_{m}^{\text{C}}$ with label $1$.
  • Unbiased samples: Interactions from $\mathcal{D}_{m}^{\text{B}}$, where responses are balanced and do not exhibit strong bias. These are considered neutral.
  • Bias-conflicting samples: Interactions that go against the dominant attribute pattern and are typically hard to learn. This includes interactions from $\mathcal{D}_{m}^{\text{A}}$ with label $1$, and from $\mathcal{D}_{m}^{\text{C}}$ with label 0.
In CAT, the examinee’s true proficiency $\theta_{i}^{*}$ is assumed constant, and traditional methods aim to make the final estimate $\theta_{i}^{T}$ closely approximate it. This work instead focuses on mitigating selection bias in the data-driven Question Selection Module to produce accurate, unbiased proficiency estimates and maintain robustness under distributional shifts.
4. Method
This section presents our approach to mitigating bias in data-driven CAT. The method addresses selection bias in the meta set through a selective mixup strategy guided by examinee attributes. By generating augmented samples from balanced response profiles, it enriches underrepresented cases and helps the Question Selection Module learn smoother, less biased decision boundaries. The approach is simple, effective, and can be readily integrated into data-driven question selection in CAT.
4.1. Overview
Our method mitigates selection bias in CAT through a two-stage framework: Cross-Attribute Examinee Retrieval and Selective Mixup Regularization, as illustrated in Figure 3. In the first stage, for each examinee in $\mathcal{D}_{m}^{\text{A}}$ and $\mathcal{D}_{m}^{\text{C}}$, the most similar counterpart is retrieved from $\mathcal{D}_{m}^{\text{B}}$ based on response patterns. In the second stage, the focus is on bias-conflicting samples, such as correct responses from $\mathcal{D}_{m}^{\text{A}}$ or incorrect responses from $\mathcal{D}_{m}^{\text{C}}$. To enrich these underrepresented cases, each examinee from $\mathcal{D}_{m}^{\text{B}}$ is paired with a counterpart from $\mathcal{D}_{m}^{\text{A}}$ or $\mathcal{D}_{m}^{\text{C}}$, and mixup is applied only when their interaction labels are identical. The final training objective combines the empirical loss on the original data with additional regularization from synthetic samples, thereby improving the robustness and generalization capability of the Question Selection Module under biased meta distributions.
Figure 3. An overview of our debiasing framework for Computerized Adaptive Testing. The method consists of two stages: (1) Cross-Attribute Examinee Retrieval, where we retrieve midperformers from $\mathcal{D}_{m}^{\text{B}}$ as neutral references for examinees in $\mathcal{D}_{m}^{\text{A}}$ and $\mathcal{D}_{m}^{\text{C}}$ based on proficiency similarity; and (2) Selective Mixup-based Regularization, where mixup is applied between retrieved midperformers and their matched counterparts (with the same label) to generate synthetic training samples. This augmentation improves the robustness of the Question Selection Module by reducing reliance on bias-aligned patterns.
4.2. Cross-Attribute Examinee Retrieval
In real-world CAT scenarios, the training data often exhibits significant selection bias: examinees in $\mathcal{D}_{m}^{\text{A}}$ are more likely to respond incorrectly, while those in $\mathcal{D}_{m}^{\text{C}}$ tend to answer most questions correctly. As a result, the training set is dominated by bias-aligned interactions, such as incorrect responses from $\mathcal{D}_{m}^{\text{A}}$ and correct responses from $\mathcal{D}_{m}^{\text{C}}$, which are easy for the Question Selection Module to fit but may lead to shortcut learning and reduced robustness. In contrast, bias-conflicting samples, including correct responses from $\mathcal{D}_{m}^{\text{A}}$ and incorrect responses from $\mathcal{D}_{m}^{\text{C}}$, are much rarer and more challenging to learn, yet they are essential for capturing deeper examinee–question relationships and ensuring fairness and generalization. To alleviate this imbalance, we leverage examinees from $\mathcal{D}_{m}^{\text{B}}$, whose balanced response patterns resemble an unbiased distribution and provide reliable references for augmenting the underrepresented bias-conflicting cases, which naturally motivates our selective mixup design.
Concretely, for each examinee $e_{a}$ with attribute $a\in\{\text{A},\text{C}\}$, we retrieve the most similar examinee $e_{b}$ with attribute $b\in\text{B}$ based on their cognitive proficiency vectors $\boldsymbol{\theta}$, estimated by the Cognitive Diagnosis Module. We compute the similarity via the $L_{2}$ norm:
(9) $\text{similarity}_{i,j}=\left\|\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{j}\right\|_{2},$
where $\boldsymbol{\theta}_{i}$ and $\boldsymbol{\theta}_{j}$ represent the latent knowledge proficiency vectors of examinees $i$ and $j$, respectively.
This retrieval process serves two purposes: first, it anchors bias-conflicting interactions with semantically similar yet statistically more balanced samples; second, it establishes the basis for subsequent augmentation. By aligning examinees in $\mathcal{D}_{m}^{\text{A}}$ and $\mathcal{D}_{m}^{\text{C}}$ with appropriate references from $\mathcal{D}_{m}^{\text{B}}$, the method effectively bridges biased extremes and enables the model to learn smoother decision boundaries in the next stage.
4.3. Selective Mixup-based Regularization
To further mitigate selection bias, we propose a selective mixup augmentation strategy (zhang2018mixup). Unlike conventional training that heavily depends on naturally occurring samples, mixup introduces synthetic interactions by interpolating instances, which substantially increases the diversity of bias-conflicting data. This is particularly important in CAT, where the scarcity of such cases otherwise causes the Question Selection Module to overfit to bias-aligned patterns and overlook minority interactions. Mixup has several key advantages in this context: it smooths decision boundaries, reduces overfitting, and improves generalization when the test distribution differs from the training distribution. Moreover, by diluting the dominance of spurious correlations embedded in the original data, mixup encourages the model to learn more balanced and robust representations. Through this targeted augmentation, the Question Selection Module is guided toward more reliable and unbiased predictions, ultimately enhancing its robustness and generalization under distribution shifts and ensuring more consistent performance across diverse examinees.
Bias-Conflicting Sample Generation
Given $q_{i}\in{\mathcal{D}_{m}^{\text{A}},\mathcal{D}_{m}^{\text{C}}}$, its most similar counterpart in terms of proficiency is first identified from $\mathcal{D}_{m}^{\text{B}}$ in the previous stage. From this matched examinee in $\mathcal{D}_{m}^{\text{B}}$, an interaction $q_{j}$ with the same true label $y$ is then sampled and paired with $q_{i}$, upon which synthetic bias-conflicting samples are generated through convex combination. With such paired questions $(q_{i},q_{j})$, we further interpolate their associated parameters to construct fake questions. The representation mixup is as follows:
(10) $\boldsymbol{Q}_{i,j}^{\text{fake}}$ $=\lambda\boldsymbol{Q}_{i}+(1-\lambda)\boldsymbol{Q}_{j},$
(11) $\boldsymbol{h}_{i,j}^{\text{diff}}$ $=\lambda\boldsymbol{h}^{\text{diff}}_{i}+(1-\lambda)\boldsymbol{h}_{j}^{\text{diff}},$
(12) $h_{i,j}^{\text{disc}}$ $=\lambda h_{i}^{\text{disc}}+(1-\lambda)h_{j}^{\text{disc}}.$
Synthetic Feature Construction
The synthesized features are then fed into the CDM, and the corresponding response probability is computed as $M(q_{i,j}^{\text{fake}};\boldsymbol{\theta},\zeta_{q_{i,j}^{\text{fake}}},\zeta_{n})$. Since the mixup is performed between samples sharing the same label, the synthetic sample inherits the same label as the originals. The loss for these augmented samples is therefore defined as:
(13) $\mathcal{L}_{\text{fake}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{q_{i,j}\in\mathcal{D}_{m}}\ell\!\left(y_{i,q_{i,j}^{\text{fake}}},M\!\left(q_{i,j}^{\text{fake}};\boldsymbol{\theta}_{i}^{*},\zeta_{q_{i,j}^{\text{fake}}},\zeta_{n}\right)\right).$
Debiasing Optimization
The final outer-level optimization objective can be rewritten as the combination of empirical risk minimization and synthetic sample regularization:
(14) $\mathcal{L}_{\text{emp}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{q_{i,j}\in\mathcal{D}_{m}}\ell\!\left(y_{i,q_{i,j}},M\!\left(q_{i,j};\boldsymbol{\theta}_{i}^{*},\zeta_{q_{i,j}},\zeta_{n}\right)\right),$
(15) $\min_{\zeta_{s}}\mathcal{L}_{\text{final}}=\mathcal{L}_{\text{emp}}+\omega\mathcal{L}_{\text{fake}}.$
where $\omega$ controls the trade-off between real and synthetic samples.
The dual-objective design encourages the model to maintain strong predictive performance on authentic data while simultaneously enhancing its robustness to bias-conflicting cases. The outer-level optimization updates the Question Selection Module parameters $\zeta_{s}$ via gradient descent on the combined loss function. A full overview of the training process, including mixup-based augmentation and bias-aware optimization steps, is provided in Algorithm 1, which outlines how the proposed strategy is integrated into the outer-loop optimization to enhance robustness and mitigate selection bias.
Algorithm 1 Debiasing in CAT Training Process
1: Initialize parameters $\boldsymbol{\theta}$ for the Cognitive Diagnosis Module using pre-trained parameters $(\zeta_{q},\zeta_{n})$, and initialize $\zeta_{s}$ for the Question Selection Module. Set learning rates $\alpha$, $\beta$, number of gradient descent steps $K$ for inner-level optimization, and number of test steps $T$.
2:while not converged do
3:  Randomly sample a mini-batch of examinees $\mathcal{D}$ with support and meta sets $\{\mathcal{D}_{u},\mathcal{D}_{m}\}$, where $\mathcal{D}_{m}=\mathcal{D}_{m}^{\text{A}}\cup\mathcal{D}_{m}^{\text{B}}\cup\mathcal{D}_{m}^{\text{C}}$.
4:  for $t=1$ to $T$ do
5:   Represent $s_{i,t}$ using the examinee’s response history.
6:   Select question $q_{i,t}$ for each examinee based on current state $s_{i,t}$ (Eq. (6)).
7:   Inner-level update: update $\boldsymbol{\theta}_{i}$ by gradient descent with learning rate $\alpha$ on the prediction loss in Eq. (7), repeated for $K$ steps, to obtain $\boldsymbol{\theta}_{i}^{*}$.
8:   For examinees in $\mathcal{D}_{m}^{\text{A}}$ and $\mathcal{D}_{m}^{\text{C}}$, retrieve the most similar examinees from $\mathcal{D}_{m}^{\text{B}}$ using Eq. (9).
9:   Apply selective mixup augmentation using Eqs. (10) and (13).
10:   Outer-level update: $\zeta_{s}\leftarrow\zeta_{s}-\beta\nabla_{\zeta_{s}}\mathcal{L}_{\text{final}}$ (Eq. (14)).
11:  end for
12:end while
13:return Estimated proficiencies $\{\boldsymbol{\theta}_{i}^{*}\}_{i=1}^{N}$ and debiased Question Selection Module $\zeta_{s}$.
Rationale and Impact
By interpolating between samples from attribute A or C and their corresponding references from attribute B, our approach introduces balanced, label-consistent synthetic data into the training process. This prevents the Question Selection Module from overfitting to dominant response patterns and encourages it to capture more nuanced relationships between question difficulty and proficiency. In effect, the model becomes less reliant on biased historical distributions, improving fairness and robustness across diverse examinees.
5. Experiments
In this section, we first introduce the dataset and experimental setup. Then, we conduct extensive experiments on real-world datasets. The code in our experiments are available at: https://anonymous.4open.science/r/MDCAT.
5.1. Experimental Settings
In this section, we introduce the datasets, the selected baselines, and the application of CAT.
DataSet Description. We conducted experiments on two widely used educational datasets. The ASSIST0910 dataset [^2] https://sites.google.com/site/assistmentsdata/home/assistment2009-2010-data contains student practice logs from the ASSISTments online tutoring system, focusing on mathematics and knowledge concept problems. The NIPS-EDU dataset [^3] https://eedi.com/projects/neurips-education-challenge, released as part of the NeurIPS 2020 Education Challenge, includes student responses collected from the Eedi learning platform. We used the interaction data from Task 3 and Task 4. For the ASSIST0910 dataset, we first removed records containing missing values and subsequently excluded examinees with fewer than 40 interactions. The dataset statistics, including the number of examinees, questions, concepts, and interactions, are presented in Table 1.
Table 1. The statistics of the datasets.
Dataset ASSIST0910 NIPS-EDU
Examinees 1, 360 4,918
questions 17,372 948
Concepts 119 86
Interactions 241,156 1,382,727
Partition. We split the students into training (60%), validation (20%), and testing (20%) sets. For evaluation, we conduct 5 rounds of validation on each dataset. In each round, we further partition the questions answered by each student into a support set ($\mathcal{D}_{u}$, 80%) and a meta set ($\mathcal{D}_{m}$, 20%). To prevent overfitting, these partitions are randomly regenerated at the beginning of each training epoch.
Evaluation.To assess model performance, we evaluate all methods based on their ability to predict binary-valued responses in the meta set $\mathcal{D}_{m}$, using two metrics: the worst-group accuracy (Worst) and the average accuracy across all samples (Avg). All models are implemented in PyTorch and trained on an NVIDIA 3090 GPU.
Baseline Methods. For Cognitive Diagnosis Modules, we adopt the 1PL IRT model (embretson2013irt) and NCDM (wang2020ncd), both of which are implemented in the EduStudio [^4] https://edustudio.ai/ framework (wu2025edustudio). For CAT frameworks, we focus on representative data-driven approaches that formulate CAT as a bi-level optimization problem and solve it via meta-learning. Specifically, we include BOBCAT (ghosh2021bobcat) and UATS (yu2024ucats), both of which jointly optimize the Question Selection Module and the student model. BOBCAT formulates the problem as a bi-level optimization task with a learnable Question Selection Module, while UATS employs a differentiable hierarchical optimization framework to enhance both efficiency and adaptability in question selection. To better analyze bias in Question Selection Module, we freeze the question-related parameters of the Cognitive Diagnosis Module during outer-loop optimization and instead initialize them with pre-trained values.
To validate the effectiveness of our method, we compare it with the following debiasing baselines, all implemented within the same bi-level training framework and adapted to the context of bias mitigation in CAT.
ERM (Empirical Risk Minimization) minimizes the average prediction loss over the entire meta set, without explicitly addressing distributional bias. It serves as a standard baseline but tends to overfit to dominant bias-aligned patterns, failing to generalize to bias-conflicting samples.
IRM (arjovsky2019irm) (Invariant Risk Minimization) encourages the model to learn feature representations that support consistent predictions across different environments. In our setting, we simulate multiple training environments by sampling examinees with different attribute values (i.e., A, B, C) and ensure that the Question Selection Module performs well across these diverse distributions.
GroupDRO (sagawa2019groupDRO) (Group Distributionally Robust Optimization) focuses on improving the worst-case performance across predefined groups. We optimize the Question Selection Module by minimizing the maximum group-wise loss, thus preventing performance collapse on underrepresented or difficult subpopulations.
Reweight (japkowicz2000reweight) directly addresses imbalance by assigning larger weights to bias-conflicting samples during training. This strategy guides the model to better capture underrepresented patterns and reduces reliance on shortcut signals in bias-aligned data.
Parameter Settings. In NCDM, 128 and 64 denote the number of neurons in the two hidden layers, which control the capacity and complexity of the network. For the $\alpha$ parameter in the Beta distribution, we set the scale to $\{0.2,0.4,0.6,0.8,1.0\}$. Similarly, for the mixing coefficient $\omega$ in the mixup loss, we use values from the same set: $\{0.2,0.4,0.6,0.8,1.0\}$. To approximate each student’s ability, we perform inner optimization by updating the student’s latent ability vector $K$ times, with $K=5$ in our experiments.
Table 2. Performance of Different CAT Models under IRT-based and NCDM-based Cognitive Diagnosis with Various Debiasing Methods on the Unbiased ASSIST0910 and NIPS-EDU Datasets. Here, Worst denotes worst‑group accuracy and Avg. denotes overall average accuracy.
Dataset ASSIST0910 NIPS-EDU
CDM-CAT IRT-BOBCAT IRT-UATS NCDM-BOBCAT NCDM-UATS IRT-BOBCAT IRT-UATS NCDM-BOBCAT NCDM-UATS
Metrics@10 Worst Avg. Worst Avg. Worst Avg. Worst Avg. Worst Avg. Worst Avg. Worst Avg. Worst Avg.
ERM 0.4088 0.6147 0.2885 0.5904 0.3206 0.5970 0.3188 0.5965 0.3383 0.6191 0.4020 0.6254 0.2847 0.6263 0.3512 0.6399
IRM 0.3890 0.6136 0.2491 0.5847 0.3277 0.5994 0.3079 0.5920 0.3245 0.6224 0.3532 0.6333 0.3056 0.6279 0.3704 0.6392
GroupDRO 0.4370 0.6285 0.3424 0.6011 0.3600 0.6003 0.2796 0.5875 0.2962 0.6253 0.3402 0.6340 0.2577 0.6247 0.3363 0.6424
Reweight 0.3625 0.6299 0.3356 0.6050 0.4551 0.6102 0.4530 0.6091 0.3133 0.6299 0.4496 0.6243 0.3446 0.6374 0.3458 0.6416
Ours 0.4489 0.6320 0.4160 0.6149 0.4426 0.6117 0.4667 0.6103 0.4489 0.6312 0.4755 0.6280 0.3520 0.6381 0.3747 0.6421
5.2. Performance in OOD Settings
Motivation. In most existing data-driven CAT studies, the meta set used to evaluate question selection performance is typically sampled directly from the examinee’s original interaction log. This procedure implicitly assumes that the observed response distribution accurately represents the examinee’s true question exposure or proficiency profile. However, in practice, this assumption is problematic. Since CAT adaptively selects questions based on estimated ability, examinees are only exposed to a small, biased subset of the entire question pool. Consequently, the meta set may reflect these selection-induced patterns rather than the examinee’s actual proficiency, leading to a biased and unrepresentative evaluation.
To better evaluate the effectiveness of our debiasing strategy in mitigating selection bias, we design an Out-of-Distribution (OOD) test setting. Unlike the conventional IID setup where the meta set is randomly sampled from the same distribution as the support set, our OOD protocol deliberately alters this distributional assumption. Specifically, for each examinee, we construct an unbiased meta set by enforcing a balanced ratio of correct and incorrect responses (1:1). The remaining interactions are then assigned to the support set, which naturally leads to a distribution shift between the two sets. Examinees whose data cannot satisfy this balancing constraint are excluded. In this way, we preserve the support–meta framework while ensuring that the evaluation truly reflects performance under distributional shifts. This controlled design provides a more rigorous and interpretable benchmark, making the OOD results highly indicative of a model’s robustness and generalization ability.
Our proposed method demonstrates strong performance across different settings, consistently surpassing baseline approaches on the Worst metric and outperforming most baselines on the Avg. metric across two distinct Cognitive Diagnostic Modules and two widely used public datasets. Specifically, on the ASSIST0910 dataset, our method achieves substantial relative improvements ranging from 2.72% to 21.5% on the Worst metric, and from 0.20% to 1.64% on the Avg. metric. Similarly, on the NIPS-EDU dataset, we observe performance gains of 1.16% to 32.69% on the Worst metric, and 0.11% to 0.21% on the Avg. metric, compared to the state-of-the-art (SOTA) results. Notably, the larger improvement margins under the IRT setting, as opposed to the NCDM setting, suggest that our method better complements the limitations of traditional models like IRT, thereby further highlighting the strengths of neural Cognitive Diagnostic Modules such as NCDM in mitigating individual performance variance. These results suggest that our approach not only provides reliable gains over ERM but also outperforms existing debiasing strategies, especially in the Worst metric, which is critical for robustness in adaptive testing.
5.3. Performance in IID Settings
As shown in Table 2 and Table 3, debiasing inevitably introduces a trade-off between IID and OOD performance. While methods such as Reweight and our approach are effective in mitigating bias and improving OOD generalization, they cannot surpass ERM or GroupDRO under IID settings, which are naturally more favorable to models optimized without debiasing constraints. Specifically, IRM achieves the best average performance on ASSIST2009 (0.6939), while ERM achieves the highest average on NIPS-EDU (0.6930). GroupDRO also maintains relatively strong IID results across both datasets. In contrast, our method focuses on improving robustness: it substantially outperforms the Reweight baseline in both Worst and Avg. accuracy under OOD conditions, while remaining competitive in-distribution. This confirms that debiasing methods, including ours, sacrifice a small amount of IID accuracy in exchange for much stronger OOD generalization, underscoring the trade-off inherent in fairness- and robustness-oriented learning.
Table 3. IID results are included as a reference to illustrate performance under in-distribution conditions.
Dataset ASSIST0910 NIPS-EDU
CDM-CAT IRT-BOBCAT IRT-BOBCAT
Metrics@10 Worst Avg. Worst Avg.
ERM 0.4609 0.6895 0.3937 0.6930
IRM 0.4248 0.6939 0.3854 0.6928
GroupDRO 0.5104 0.6630 0.3535 0.6922
Reweight 0.4179 0.6155 0.3576 0.6622
Ours 0.5310 0.6599 0.4911 0.6695
5.4. Hyperparameter Analysis
We conduct a hyperparameter analysis of the mixup regularization weight $\omega$ on the NIPS-EDU dataset, evaluating its impact on Worst accuracy across two Cognitive Diagnosis Models and two CAT frameworks. As shown in Figure 4, most CDM-CAT combinations achieve their best performance when $\omega$ is set to 0.8 or 1.0. The overall trend follows a rise-then-fall pattern: performance initially improves as $\omega$ increases, peaking around 0.8 or 1.0, and then gradually declines beyond that point.Notably, our method demonstrates relatively stable performance across a wide range of $\omega$ values. This suggests that the model is not overly sensitive to the specific setting of the regularization strength, which enhances its practical applicability. Even when $\omega$ deviates slightly from the optimal value, the model maintains strong worst-group accuracy, indicating robust debiasing behavior and consistent generalization across configurations.
Figure 4. The effect of the regularization weight $\omega$ on Worst accuracy shows that performance peaks around 0.2–1.0 and remains relatively stable across a wide range.
5.5. Ablation Study
We conduct ablation studies to examine three mixup strategies: Mixup-Self, Mixup-Inner, and Mixup-B. In Mixup-Self, interactions are interpolated within the same examinee, which introduces synthetic variation into an individual learner’s profile and can occasionally improve robustness. Mixup-Inner interpolates the most similar examinees within the same attribute, thereby smoothing local representations and reducing overfitting to individual-specific patterns. In contrast, Mixup-B interpolates biased examinees (e.g., from Attribute A or C) with balanced counterparts (Attribute B). By explicitly synthesizing bias-conflicting interactions, this strategy enriches underrepresented patterns and more effectively mitigates selection bias. As shown in Figure 5, Mixup-B consistently achieves superior performance, particularly on the large-scale ASSIST0910 dataset, where the greater number of questions provides higher diversity and amplifies the benefits of cross-attribute mixing. These results highlight both the effectiveness of cross-attribute interpolation for debiasing and the potential of self-interaction strategies, which, although less stable, remain a promising direction for future exploration.
(a) Ablation results of different mixup strategies on ASSIST0910.
(b) Ablation results of different mixup strategies on NIPS-EDU.
Figure 5. Ablation results of different mixup strategies (Mixup-Self, Mixup-Inner, and Mixup-B) on ASSIST0910 and NIPS-EDU in terms of Worst@10 and Avg.@10.
5.6. Mitigating Distribution Shift in Question Selection
While previous results have demonstrated the effectiveness of our method in improving Question Selection Module performance under OOD conditions, we further investigate the underlying mechanism by analyzing the distributional differences introduced by our approach. Specifically, Figure 6 illustrates the distribution of correct-response ratios among examinees under both the ERM baseline and our proposed method. For examinees with a lower overall correct-response ratio, our method significantly shifts the distribution toward a more moderate range, concentrating around 0.4. This suggests that fewer overly difficult questions are being selected, thereby reducing over-challenging effects. Conversely, for high-performing examinees, our method attenuates the skewness toward extremely easy questions by centering the distribution around 0.6, rather than approaching 1.0 as observed in ERM. This indicates a reduction in overfitting to bias-aligned patterns.
These results show that our method effectively mitigates the distribution shift between selected questions and the meta set, leading to a more balanced and representative question selection across diverse examinees. By increasing the diversity of bias-conflicting samples, our method adjusts the selected-question distribution to better preserve individual examinee characteristics while simultaneously aligning it more closely with a fair target distribution. This ensures that question selection reflects true proficiency rather than reinforcing spurious patterns from imbalanced training data. Our approach enhances generalization and supports more reliable proficiency estimation.
Figure 6. Distribution of Selected-Question Correct Ratios Across Examinees.
6. Conclusion
In this study, we identify and address the selection bias problem in data-driven Computerized Adaptive Testing, particularly the bias arising from meta-set distribution imbalance during training. We propose a mixup strategy guided by moderately performing students to augment underrepresented interactions and improve the generalization of the Question Selection Module. This method is simple, annotation-free, and can be seamlessly integrated into standard CAT training pipelines. Experiments on two benchmark datasets show that our approach significantly enhances both fairness and accuracy. By enhancing the diversity of bias-conflicting samples through synthetic data generation, the model learns more robust and less biased decision boundaries.This work highlights the importance of fairness in intelligent educational systems and suggests correcting bias directly from the selection process as a promising future direction.
7. GenAI Usage Disclosure
This paper involved the use of generative AI tools (e.g., ChatGPT) for non-substantive writing assistance. Specifically, AI was used to help with language polishing, improving clarity and grammar in some parts of the paper. In addition, AI-assisted tools were employed to generate code for visualizing experimental results, as well as to provide suggestions on the overall writing structure and organization. All technical content, experimental design, and core contributions were conceived and implemented by the authors.
Acknowledgements.
This research was partially supported by grants from National Science and Technology Major Project under Grant (No. 2021ZD0111802), the National Natural Science Foundation of China (No. 62376086, 62406096), and the Fundamental Research Funds for the Central Universities of China (No. JZ2025HGTG0289). the China Postdoctoral Science Foundation (under Grant 2024M760722). The computation is completed on the HPC Platform of Hefei University of Technology.